# High Fidelity Text-Guided Music Generation and Editing via Single-Stage Flow Matching

### v1

@@ Abstract
We introduce a simple and efficient text-controllable high-fidelity music generation and editing model. It operates on sequences of continuous latent representations from a low frame rate 48 kHz stereo variational auto encoder codec that eliminates the information loss drawback of discrete representations. Based on a diffusion transformer architecture trained on a flow-matching objective the model can generate and edit diverse high quality stereo samples of variable duration, with simple text descriptions. We also explore a new regularized latent inversion method for zero-shot test-time text-guided editing and demonstrate its superior performance over naive denoising diffusion implicit model (DDIM) inversion for variety of music editing prompts. Evaluations are conducted on both objective and subjective metrics and demonstrate that the proposed model is not only competitive to the evaluated baselines on a standard text-to-music benchmark - quality and efficiency-wise - but also outperforms previous state of the art for music editing when combined with our proposed latent inversion. Samples are available at https://melodyflow.github.io.
@@ 1 Introduction
Text-conditioned music generation has made tremendous progress in the past two years (Schneider et al., 2023; Huang et al., 2023; Agostinelli et al., 2023; Copet et al., 2024; Ziv et al., 2023; Liu et al., 2023; Li et al., 2023). The dominant approach involves representing audio as a sequence of compressed discrete or continuous representations and training a generative model on top of it. Two dominant generative model architectures have emerged, one based on Language Models (LMs) (Agostinelli et al., 2023; Copet et al., 2024), the other on diffusion (Schneider et al., 2023; Huang et al., 2023; Liu et al., 2023; Li et al., 2023). A third method sometimes referred to as discrete diffusion relies on non-autoregressive masked token prediction (Ziv et al., 2023; Garcia et al., 2023). The target level of audio fidelity depends on the models and some have already successfully generated 44.1 kHz or high stereo signals (Schneider et al., 2023; Li et al., 2023; Evans et al., 2024a). However no such work has really discussed the implications of aiming for higher fidelity or stereo and what performance compromise it implies from the latent representation and generation model perspective.
@@
Due to the growing popularity of diffusion models in computer vision, a new area of research has emerged focusing on text-controlled audio editing (Wang et al., 2023; Lin et al., 2024; Garcia et al., 2023; Wu et al., 2023; Novack et al., 2024; Zhang et al., 2024; Manor and Michaeli, 2024). The creative process for sound design typically involves multiple iterations and using efficient editing methods is an effective approach to achieving this. Music editing constitutes an open ended list of tasks such as inpainting/outpainting, looping, instrument or genre swapping, vocals removal, lyrics editing, tempo control or recording conditions modification (e.g. from studio quality to a concert setting). Some of those tasks have been addressed in recent works with specialised models (Wang et al., 2023; Garcia et al., 2023; Lin et al., 2024; Wu et al., 2023; Copet et al., 2024) or zero-shot editing methods from the computer vision domain that are exclusive to diffusion models (Novack et al., 2024; Zhang et al., 2024; Manor and Michaeli, 2024). However, none of these have demonstrated high fidelity generic style transfer capabilities, due to either lack of high quality data or foundational music generation model, or design choices that do not generalise to any kind of editing task. The question of inference speed is key for creatives and the music domain is particularly challenging due to the high fidelity (48 kHz stereo) requirement in the sound design process. Recently the Flow Matching (FM) generative modeling formulation has been introduced (Lipman et al., 2022), that consists in building optimal transport paths between data and noise samples. It is a more robust and stable alternative for training diffusion models with notably faster inference and was successfully applied to train foundational speech (Le et al., 2024) and audio (Vyas et al., 2023) generative models.
@@
In this work we present MelodyFlow, a single-stage text-conditioned FM model for instrumental music generation and editing. The model operates on continuous representations of a low frame rate Variational Audio Encoder (VAE) codec, which is crucial for handling high-fidelity stereo generation efficiently compared to models operating on discrete neural codecs. Our evaluations show that the proposed model is competitive with evaluated baselines on text-to-music generation while being remarkably efficient, with a latency of less than 10 seconds for a 30 seconds sample generation (64 inference steps). Additionally, thanks to the versatility of FM, MelodyFlow is compatible with any zero-shot test-time editing method such as DDIM inversion (Song et al., 2020). We enhance the editability of the FM inversion process with a regularized latent inversion method inspired by that of (Garibi et al., 2024) for diffusion-based image editing. Both our objective and subjective evaluations on music editing indicate that MelodyFlow can support a diversity of editing tasks on real songs without any finetuning, achieving fast music editing with remarkable consistency, text-adherence and minimal quality degradation compared with original samples. In addition we conduct an ablation study on the importance of the key design choices on the overall model quality/efficiency trade off.
@@
Our contributions: (i) We introduce an efficient single-stage text-to-music FM model to generate 48 kHz stereo samples of up to 30 seconds, with enhancements in both the audio latent representation and generative model, striking a better balance between quality and efficiency. (ii) We present a novel regularized FM inversion method capable of performing faithful zero-shot test-time text-guided editing on various axes while maintaining coherence with the original sample. (iii) We conduct comprehensive objective and human evaluations on the key design choices underlying our approach.
@@ 2 Method
MelodyFlow is a non-causal transformer-based flow matching model conditioned on T5 text representations. It operates on sequences of latent audio representations from a neural quantizer-free audio codec with a VAE bottleneck. Combined with latent inversion, the model is able to perform text-guided editing of real or generated audio samples.
@@ 2.1 Latent audio representation
Our audio codec derives from EnCodec (Défossez et al., 2022) with additional features from the Descript Audio Codec (DAC) (Kumar et al., 2024) (snake activations, band-wise STFT discriminators) and (Evans et al., 2024a) (VAE bottleneck, perceptual weighting). A convolutional auto-encoder encodes the raw waveform into a sequence of latent bottleneck representations, its frame rate function of the convolution strides. Audio fidelity is enforced by multi-scale STFT reconstruction losses (both genuine and adversarial). The VAE bottleneck enforces smoothness in the latent representation space, as opposed to discretization bottlenecks such as Residual Vector Quantization (RVQ) (Défossez et al., 2022). Although (Vyas et al., 2023) have successfully trained a flow matching model on continuous latent representations extracted from the encoder of a discrete audio codec (before the RVQ layer), we argue that the latent space of such a codec is not the best suited for continuous modeling. Moreover while discrete representations have shown promising results on generating 32 kHz mono music data (Copet et al., 2024), adapting such models for higher fidelity stereo generation requires either scaling the vocabulary size, frame rate or number of residual projections to a point where discrete autoregressive LM becomes prohibitively more expensive than non-autoregressive continuous FM. Indeed single-stage LMs require some degree of parallel token prediction to keep the inference time acceptable (Le Lan et al., 2024). We show in section 4.3.1 that a VAE quantizer-free codec can operate at a much lower frame rate than its discrete equivalent for a better level of fidelity, reducing the cost of generation model training and inference by essentially the same factor.
@@ 2.2 Conditional flow matching model
Given an audio sample $\mathbf{a}\in\mathbb{R}^{D\times f_{s}}$, a sequence $\mathbf{x}\in\mathbb{R}^{L\times d}$ of latent representations is extracted by the neural codec. Flow matching models optimal transport paths that map a sequence $\mathbf{\epsilon}\in\mathbb{R}^{L\times d}\sim\mathcal{N}(0,I)$ to $\mathbf{x}$ trough a linear transformation - function of the flow step $t$ following equation 1.
$\mathbf{z}_{t}=t\mathbf{x}+(1-t)\mathbf{\epsilon},t\in[0,1]$ (1)
During training, $t$ is randomly sampled and the neural network $\Theta$ is trained to estimate the derivative $d\mathbf{z}_{t}/dt$ conditioned on $t$ and a text description $c$.
$d\mathbf{z}_{t}/dt=v_{\Theta}(\mathbf{z}_{t},t,c)=\mathbf{x}-\mathbf{\epsilon}$ (2)
@@
$\mathbf{x}_{generated}=\textbf{ODE}_{0\xrightarrow{}1}(\epsilon,c)$ (3)
[…]
@@ 2.3 Text-guided editing through latent inversion
Due to the bijective nature of the FM formulation (where given a text condition, each latent sequence is mapped to a single noise vector), the model is compatible with existing latent inversion methods such as DDIM inversion (Song et al., 2020). Given the latent representation $\mathbf{x}_{orig}$ of an existing audio with an optional accompanying caption $c\in\{\varnothing,c_{orig}\}$, the model can serve to estimate its corresponding noise (or intermediate) representation $\mathbf{z}_{t_{edit}}=\textbf{ODE}_{t_{edit}\xleftarrow{}1}(\mathbf{x}_{orig},c)$ by running the ODE solver in the backward direction until an intermediary time step $t_{edit}$. It is important to note that neither $\mathbf{x}_{orig}$ nor $c_{orig}$ have to be human-generated. $\mathbf{x}_{orig}$ could be the output of a preceding text-to-music or music editing process and $c_{orig}$ could result from a music captioning model. Given the intermediary representation $\mathbf{z}_{t_{edit}}$, the ODE forward process can be conditioned on a new text description $c_{edit}$ that materialises the editing prompt: $\mathbf{x}_{edit}=\textbf{ODE}_{t_{edit}\xrightarrow{}1}(\mathbf{z}_{t_{edit}},c_{edit})$. A good inversion process should accurately reconstruct the input when $c_{edit}=c_{orig}$, as shown in equation 4.
$\mathbf{x}_{edit}=\textbf{ODE}_{t_{edit}\xrightarrow{}1}(\textbf{ODE}_{t_{edit}\xleftarrow{}1}(\mathbf{x}_{orig},c\in\{\varnothing,c_{orig}\}),c_{orig})\approx\mathbf{x}_{orig}$ (4)
@@
In such case when swapping $c_{orig}$ for $c_{edit}$ in the $t_{edit}\xrightarrow{}1$ forward direction, the expectation is for the generated audio to preserve some consistency with the original while being faithful to the editing prompt. However in practice it was observed by (Mokady et al., 2023) that DDIM inversion suffers from poor editability due to the classifier free guidance.
2.4 Optimizing for editing
The quality of music editing depends on the codec, FM model and the latent inversion method, while we know speed is a crucial factor in the sound design creative process. Some of our design choices are critical in striking a good balance between quality and efficiency.
2.4.1 Codec VAE bottleneck
We show in section 4.3.1 that swapping the RVQ by a VAE bottleneck is key to increase the generation model efficiency and the quality of its outputs. Namely this allows better waveform reconstruction performance for a much lower frame rate, which unlocks faster inference and scaling to high-fidelity stereo audio (shown in section 4.3.2).
2.4.2 Minibatch coupling
(Tong et al., 2023; Pooladian et al., 2023) expanded over prior work on flow matching modeling by sampling pairs $(\mathbf{x},\mathbf{\epsilon})$ from the joint distribution given the by the optimal transport plan between the data $\mathbf{X}=\{{\mathbf{x}^{(i)}}\}_{i=1}^{B}$ and noise $\mathbf{E}=\{{\mathbf{\epsilon}^{(i)}}\}_{i=1}^{B}$ samples within a batch of size $B$. Essentially this translates into running the Hungarian algorithm so as to find the permutation matrix $\mathbf{P}$ that minimizes $||\mathbf{X}-\mathbf{P}\mathbf{E}||_{2}^{2}$. They demonstrate it results in straighter optimal transport paths during inference (that are closer to the theoretical linear mapping assumption between noise and data samples) and consequently offers better quality-efficiency trade offs. We shed light on the importance of mini-batch coupling in section 4.3.3 and we underline the overall benefit of the FM model design choices on its efficiency in appendix A.2.2.
2.4.3 Regularized latent inversion
(Garibi et al., 2024) recently proposed ReNoise, an iterative renoising mechanism applied at each inversion sampling step of a diffusion model. It enables building reversible diffusion trajectories with a good reconstruction/editability balance thanks to a noise regularization applied at each inversion step. We propose to follow a similar reasoning but in the flow matching setting. The theoretical grounds of flow matching are about estimating straight trajectories between noise and data samples, which helps in achieving efficient inference. However in practice those trajectories are never completely straight and naive DDIM inversion suffers from two problems.
@@ 2.4.3 Regularized latent inversion
    When running the ODE solver, any pair of successive $(\mathbf{z}_{t_{1}},\mathbf{z}_{t_{2}})$ along the inversion path usually has estimated velocities $v_{\Theta}(\mathbf{z}_{t_{1}},t_{1},c)\neq v_{\Theta}(\mathbf{z}_{t_{2}},t_{2},c)$. Building a fully reversible inversion path requires estimating $\mathbf{z^{\prime}}_{t_{2}}$ such that $v_{\Theta}(\mathbf{z}_{t_{1}},t_{1},c)\approx v_{\Theta}(\mathbf{z^{\prime}}_{t_{2}},t_{2},c)$, for example using the convergence property demonstrated by (Garibi et al., 2024).
@@
    The distribution of predicted velocities tends to shift away from that of training, partially due to the classifier free guidance. One way for overcome this is to regularize them via gradient descent. Given a manually constructed $\mathbf{\tilde{z}}_{t}=\mathbf{x}t+\epsilon(1-t)$ and a real $\mathbf{z}_{t}$ along the inversion trajectory, and corresponding predictions $v_{\Theta}(\mathbf{z}_{t},t,c)$ and $\tilde{v}_{\Theta}(\mathbf{\tilde{z}}_{t},t,c)$. We arrange model predictions in 4x4 patches and compute the average Kullback-Leibler (KL) divergence $\mathcal{L}_{patchKL}$ between corresponding patches.
Our adaptation of the ReNoise inversion algorithm for flow matching is detailed in algorithm 1 and we provide a comparative analysis with DDIM inversion in sections 4.3.4 and 4.3.5. The main difference between our method and ReNoise is that given the FM formulation we only enforce the model prediction regularization, but neither de-correlation nor noise correction steps. In appendix A.2.3 we compare our approach with ReNoise by reformulating the FM prediction into a noise prediction task.
@@
Sequence of audio latents $\mathbf{x}$. Number of ODE backward steps $S$. Original text description $c\in\{\varnothing,c_{orig}\}$. ReNoise iteration steps $K$ with weights $\{w_{k}\}_{k=1}^{K}$, KL regularization weight $\lambda_{KL}$.
A noisy latent $\mathbf{z}_{T_{edit}}$ such that $\textbf{ODE}_{T_{edit}\xrightarrow{}1}(\mathbf{z}_{T_{edit}},c_{orig})\approx\mathbf{x}$.
@@
              $\mathbf{\tilde{z}}^{(k-1)}_{t-\Delta{t}}\leftarrow\mathbf{x}t+\epsilon(1-t)$

### v2

# High Fidelity Text-Guided Music Editing via Single-Stage Flow Matching

@@ Abstract
We introduce MelodyFlow, an efficient text-controllable high-fidelity music generation and editing model. It operates on continuous latent representations from a low frame rate 48 kHz stereo variational auto encoder codec. Based on a diffusion transformer architecture trained on a flow-matching objective the model can edit diverse high quality stereo samples of variable duration, with simple text descriptions. We adapt the ReNoise latent inversion method to flow matching and compare it with the original implementation and naive denoising diffusion implicit model (DDIM) inversion on a variety of music editing prompts. Our results indicate that our latent inversion outperforms both ReNoise and DDIM for zero-shot test-time text-guided editing on several objective metrics. Subjective evaluations exhibit a substantial improvement over previous state of the art for music editing. Code and model weights will be publicly made available. Samples are available at https://melodyflow.github.io.
@@ 1 Introduction
Text-conditioned music generation has made tremendous progress in the past two years (Schneider et al., 2023; Huang et al., 2023; Agostinelli et al., 2023; Copet et al., 2024; Ziv et al., 2023; Liu et al., 2023b; Li et al., 2023; Prajwal et al., 2024). The prevailing method for audio representation involves compressing the waveform into a series of discrete or continuous tokens, and then training a generative model on top of those. Two dominant generative model architectures have emerged, one based on autoregressive Language Models (LMs) (Agostinelli et al., 2023; Copet et al., 2024), the other on diffusion (Schneider et al., 2023; Huang et al., 2023; Liu et al., 2023b; Li et al., 2023; Prajwal et al., 2024). A third method sometimes referred to as discrete diffusion relies on non-autoregressive masked token prediction (Ziv et al., 2023; Garcia et al., 2023). The target level of audio fidelity depends on the models and some have already successfully generated 44.1 kHz or high stereo signals (Schneider et al., 2023; Li et al., 2023; Evans et al., 2024a).
@@
The increasing popularity of diffusion models in computer vision has led to the emergence of a new area of research focused on text-controlled audio editing (Wang et al., 2023; Lin et al., 2024; Garcia et al., 2023; Wu et al., 2023; Novack et al., 2024; Zhang et al., 2024; Manor & Michaeli, 2024). The sound design process often involves multiple iterations, and using efficient editing methods is a key approach to achieving this effectively. Music editing encompasses a wide range of tasks, including but not limited to: inpainting/outpainting, looping, instrument or genre swapping, vocals removal, lyrics editing, tempo control, and recording conditions modification (e.g. from studio quality to a concert setting). Recent works have addressed some of these tasks using specialized models (Wang et al., 2023; Garcia et al., 2023; Lin et al., 2024; Wu et al., 2023; Copet et al., 2024) or zero-shot editing methods from the computer vision domain, which are exclusive to diffusion models (Novack et al., 2024; Zhang et al., 2024; Manor & Michaeli, 2024). Despite recent efforts, no approach has yet shown the ability to perform high-fidelity generic style transfer across various music editing tasks. This limitation can be attributed to several factors, including insufficient high-quality data, inadequate foundational music generation models, and design choices that fail to generalize effectively to diverse editing tasks. Inference speed is crucial for creatives, and the music domain presents a unique challenge due to the high-fidelity (48 kHz stereo) requirement in the sound design process. Recently Lipman et al. (2022) proposed the Flow Matching (FM) generative modeling formulation, which involves constructing optimal transport paths between data and noise samples. Flow Matching (FM) offers a more robust and stable approach to training diffusion models, with the added benefit of faster inference. This method has been successfully applied to train foundational speech (Le et al., 2024) and audio (Vyas et al., 2023) generative models. For the music domain Prajwal et al. (2024) utilized a two-stage FM model for text-guided music generation, where the first stage generates semantic features and the second stage generates acoustic features.
@@
In this work we present MelodyFlow, a single-stage text-conditioned FM model designed for instrumental music generation and editing. The model operates on continuous representations of a low frame rate Variational Audio Encoder (VAE) codec. Additionally, thanks to the versatility of FM, MelodyFlow is compatible with any zero-shot test-time editing method such as DDIM inversion (Song et al., 2020) or ReNoise (Garibi et al., 2024). We enhance the editability of the FM inversion by adapting the latent inversion of Garibi et al. (2024) to the FM formulation. Both our objective and subjective evaluations on music editing indicate that MelodyFlow can support a diversity of editing tasks on real songs without any finetuning, achieving fast music editing with remarkable consistency, text-adherence and minimal quality loss compared with original samples. In addition we conduct an ablation study on the importance of the key design choices on the overall model quality/efficiency trade off.
@@
Our contributions: (i) We introduce the first of its kind single-stage text-to-music FM model to generate and edit 48 kHz stereo samples of up to 30 seconds, with enhancements in both the audio latent representation and generative model, striking a better balance between quality and efficiency. (ii) We explore a novel regularized FM inversion method capable of performing faithful zero-shot test-time text-guided editing on various axes while maintaining coherence with the original sample. (iii) We publicly release the code and model weights to foster research on music editing.
@@ 2 Method
MelodyFlow combines a continuous audio codec, a text-conditioned Diffusion Transformer (DiT) FM model and a regularized latent inversion method. The model can perform text-guided editing of real or generated audio samples. The overall editing process is depicted in the Figure 1.
Figure 1: Overview of the MelodyFlow editing process. A waveform is encoded into $\mathbf{x}_{src}$ before being fed to the ODE solver. Step-by-step, the DiT predicts the velocity $\delta$ from data to noise, while being regularized against the prediction of an artificially constructed $\mathbf{\tilde{z}}_{t}$ so as to enhance editability. Once the target inversion flow step $T_{edit}$ has been reached, the model is used in the classic generation setting (bottom of the Figure, from right to left), except that the starting latent $\mathbf{z}_{t_{edit}}$ has been estimated so as to achieve better editability and consistency with the source waveform.
@@ 2.1 Latent audio representation
Our codec derives from EnCodec (Défossez et al., 2022) with additional features from the Descript Audio Codec (DAC) (Kumar et al., 2024) (snake activations, band-wise STFT discriminators) and Evans et al. (2024a) (KL-regularized bottleneck, perceptual weighting). A convolutional auto-encoder encodes the waveform into a sequence of latent bottleneck representations, its frame rate function of the convolution strides. Audio fidelity is enforced by multi-scale STFT reconstruction losses complemented by the sum and difference STFT loss for stereo support (Steinmetz et al., 2020).
@@ 2.2 Conditional flow matching model
Given an audio sample $\mathbf{a}\in\mathbb{R}^{D\times f_{s}}$, a sequence $\mathbf{x}\in\mathbb{R}^{L\times d}$ of latent representations is extracted by the neural codec. FM models the optimal transport paths that map a sequence $\mathbf{\epsilon}\in\mathbb{R}^{L\times d}\sim\mathcal{N}(0,I)$ to $\mathbf{x}$ trough a linear transformation - function of the flow step $t$ following equation 2.2.
$\mathbf{z}_{t}=t\mathbf{x}+(1-t)\mathbf{\epsilon},t\in[0,1]$
During training, $t$ is randomly sampled and the DiT $\Theta$ is trained to estimate $d\mathbf{z}_{t}/dt$ conditioned on $t$ and a text description $c$.
$d\mathbf{z}_{t}/dt=v_{\Theta}(\mathbf{z}_{t},t,c)=\mathbf{x}-\mathbf{\epsilon}$
@@
$\mathbf{x}_{generated}=\textbf{ODE}_{0\xrightarrow{}1}(\epsilon,c)$
After the latents have been estimated they are fed to the codec decoder to materialize the waveform. Kingma & Gao (2024) show that the flow step sampling density during training plays an important role in model performance. In our implementation $t$ is sampled from a logit-normal distribution (Karras et al., 2022; Esser et al., 2024).
@@ 2.3 Text-guided editing through latent inversion
Due to the bijective nature of the FM formulation (where given a text condition, each latent sequence is mapped to a single noise vector), the model is compatible with existing latent inversion methods such as DDIM inversion (Song et al., 2020). Given the latent representation $\mathbf{x}_{src}$ of an existing audio with an optional accompanying caption $c\in\{\varnothing,c_{src}\}$, the model can estimate its corresponding noise (or intermediate) representation $\mathbf{z}_{t_{edit}}=\textbf{ODE}_{t_{edit}\xleftarrow{}1}(\mathbf{x}_{src},c)$ by running the ODE solver in the backward direction until an intermediary time step $t_{edit}$ (top of the Figure 1). Given the intermediary representation $\mathbf{z}_{t_{edit}}$, the ODE forward process can be conditioned on a new text description $c_{edit}$ that materialises the editing prompt: $\mathbf{x}_{edit}=\textbf{ODE}_{t_{edit}\xrightarrow{}1}(\mathbf{z}_{t_{edit}},c_{edit})$. A good inversion process should accurately reconstruct the input when $c_{edit}=c_{src}$, as shown in equation 2.3.
$\mathbf{x}_{edit}=\textbf{ODE}_{t_{edit}\xrightarrow{}1}(\textbf{ODE}_{t_{edit}\xleftarrow{}1}(\mathbf{x}_{src},c\in\{\varnothing,c_{src}\}),c_{src})\approx\mathbf{x}_{src}$
@@
In such case when swapping $c_{src}$ for $c_{edit}$ in the $t_{edit}\xrightarrow{}1$ forward direction, the expectation is for the generated audio to preserve some consistency with the source while being faithful to the prompt. However in practice it was observed by Mokady et al. (2023) that DDIM inversion suffers from poor editability due to the classifier free guidance.
2.4 Regularized latent inversion
Even though FM consists in estimating straight trajectories, in practice those are never completely straight and the edited samples do not preserve enough consistency with the source.
@@ 2.4 Regularized latent inversion
    The distribution of predicted velocities tends to shift away from that of training due to the classifier free guidance (Mokady et al., 2023), which can lead to divergence of the inversion trajectory. This was observed by Parmar et al. (2023) with $\epsilon$-prediction, which they address by adding an autocorrelation regularization during inversion to preserve the statistical properties of the predictions.
@@
    Any pair of successive $(\mathbf{z}_{t},\mathbf{z}_{t-\Delta t})$ along the inversion path usually has estimated velocities $v_{\Theta}(\mathbf{z}_{t},t,c)\neq v_{\Theta}(\mathbf{z}_{t-\Delta t},t-\Delta t,c)$, which affects reversibility (hence the consistency with the source sample). Building a fully reversible inversion path requires estimating $\mathbf{z^{\prime}}_{t-\Delta t}$ such that $v_{\Theta}(\mathbf{z}_{t},t,c)\approx v_{\Theta}(\mathbf{z^{\prime}}_{t-\Delta t},t-\Delta t,c)$, for example following Garibi et al. (2024).
ReNoise (Garibi et al., 2024) addresses those two problems by combining both $\epsilon$-prediction regularization and reversible inversion trajectory estimation. Applying ReNoise to FM requires either (1) reformulating FM as $\epsilon$-prediction or (2) adapting the regularization mechanism. Indeed since our FM model predicts the velocity $v_{\Theta}(\mathbf{z}_{t},t,c)=\mathbf{x}-\mathbf{\epsilon}$ and ReNoise operates on noise predictions, applying ReNoise to FM (1) requires subtracting the source latent $\mathbf{x}_{src}$ from $v_{\Theta}$ to try and isolate and regularize $\mathbf{\epsilon}$ directly. However in such setting the inversion diverges when conditioning on text ($c_{src}$) and using CFG (appendix A.2.4), likely due to $v_{\Theta}(\mathbf{z}_{t},t,c)-\mathbf{x}_{src}$ not properly removing the signal component of $\mathbf{x}-\epsilon$ at lower flow steps. To prevent this behavior we propose to (2) directly regularize the FM prediction using only the KL regularization from Garibi et al. (2024). An thorough comparison between the considered approaches can be found in the sections 4.3.1, 4.3.2 and 4.5.2.
The Algorithm 1 details our proposed inversion. Each iteration consists in estimating a reversible inversion point $\mathbf{z}_{t-\Delta t}$ from a source point $\mathbf{z}_{t}$ such that $v_{\Theta}(\mathbf{z}_{t-\Delta t},t-\Delta t,c)\approx v_{\Theta}(\mathbf{z}_{t},t,c)$. In such case the jump from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-\Delta t}$ is considered reversible. This is done iteratively in $K$ steps following the convergence property of Garibi et al. (2024). During each of those steps, the model prediction is regularized against the prediction of an artifically constructed $\mathbf{\tilde{z}}_{t-\Delta{t}}$ (also shown in the Figure 1).
@@
Sequence of audio latents $\mathbf{x}$. Number of ODE backward steps $S$. Source text description $c\in\{\varnothing,c_{src}\}$. K regularization steps with weights $\{w_{k}\}_{k=1}^{K}$, KL regularization weight $\lambda_{KL}$.
A noisy latent $\mathbf{z}_{T_{edit}}$ such that $\textbf{ODE}_{T_{edit}\xrightarrow{}1}(\mathbf{z}_{T_{edit}},c_{src})\approx\mathbf{x}$.
@@
              $\mathbf{\tilde{z}}^{(k-1)}_{t-\Delta{t}}\leftarrow\mathbf{x}(t-\Delta{t})+\epsilon(1-(t-\Delta{t}))$
@@
2.5 Improving Flow Matching for text-to-music generation
2.5.1 Codec bottleneck
Recently Prajwal et al. (2024) trained a two-stage music FM model on continuous latent representations, but both the semantic and acoustic latent representations where trained with a discretization objective (HuBERT semantic features and RVQ-regularized codec). The concurrent work of Evans et al. (2024b) demonstrated long form music generation capabilities by using a KL-regularized bottleneck in their codec with a temporal downsampling as low as 21.5 Hz. However none of these works have carefully investigated the influence of the bottleneck regulariser on both music reconstruction and generation performance, all other things being equal. Indeed Rombach et al. (2022) - a seminal work on VAE for image generation - note that LDMs trained in VQ-regularized latent spaces achieve better sample quality than KL-regularized ones. Our ablation in section 4.4 leads to a different conclusion. Using a KL-regularizer achieves indeed better music reconstruction and generation performance for a much lower frame rate, which is key for faster inference.
2.5.2 Minibatch coupling
Tong et al. (2023) and Pooladian et al. (2023) expanded over prior work on FM modeling by sampling pairs $(\mathbf{x},\mathbf{\epsilon})$ from the joint distribution given the by the optimal transport plan between the data $\mathbf{X}=\{{\mathbf{x}^{(i)}}\}_{i=1}^{B}$ and noise $\mathbf{E}=\{{\mathbf{\epsilon}^{(i)}}\}_{i=1}^{B}$ samples within a batch of size $B$. Essentially this translates into running the Hungarian algorithm so as to find the permutation matrix $\mathbf{P}$ that minimizes $||\mathbf{X}-\mathbf{P}\mathbf{E}||_{2}^{2}$. They demonstrate it results in straighter optimal transport paths during inference (that are closer to the theoretical linear mapping assumption between noise and data samples) and consequently offers better quality-efficiency trade offs. We shed light on the importance of mini-batch coupling in sections 4.5.1 and 4.5.2 where we underline the overall benefit of our FM model design choices on both music generation and editing.